# ORBIT-2: Scaling Exascale Vision Foundation Models for Weather and Climate Downscaling

### v1

@@ V-E Fine-Tuning & Inference
US Regional Fine-Tuning. Following pretraining on the datasets in Table I, we fine-tune ORBIT-2 models with 9.5M and 126M parameters on two tasks: (1) US-specific downscaling of ERA5 and DAYMET data from 28 km to 7 km, evaluated against 7 km DAYMET observations for both daily total precipitation and minimum temperature; and (2) global downscaling of 28 km ERA5 precipitation to 7 km, evaluated against IMERG observations. These tasks evaluate both regional accuracy and global generalization. The current fine-tuning results are limited to 9.5M and 126M models. We will update the manuscript soon to also report fine-tuning results for the 1B and 10B parameter models.
@@
Table IV(a) summarizes results for daily minimum temperature downscaling. Both 9.5M and 126M models accurately reconstruct high-resolution temperature at 7 km resolution and are capable of capturing extremes, with the larger 126M model consistently outperforming the smaller 9.5M model across all metrics. Notably, the 126M model achieves an $R^{2}$ of 0.999 and SSIM of 0.987, establishing a new benchmark for temperature downscaling at 7 km resolution.
@@
Fig. 7(a) provides corresponding spectral analysis for the downscaling results in Table IV(a) by comparing the spatial power spectra of the two model sizes. The 126M model accurately captures high-frequency content, closely matching the high frequency spectral characteristics of the DAYMET observation ground truth. In contrast, the 9.5M model deviates from the ground truth at high frequencies. This demonstrates the larger model’s ability to resolve fine-scale spatial variability, emphasizing the value of increased model capacity for high-fidelity climate downscaling.
@@
Table IV(b) presents downscaling results for daily total precipitation—one of the most challenging variables due to its high spatial variability and localized extremes. ORBIT-2 demonstrates strong performance, closely matching observations when downscaling to 7 km resolution. Fig. 7(b) provides a visual comparison of 7 km DAYMET ground truth and ORBIT-2’s 7 km prediction. Since ERA5 (a reanalysis product) and DAYMET (observations) have different statistical properties, we combine them at 28 km resolution as the input for downscaling to 7 km. The 126M model consistently outperforms the 9.5M model across all evaluation metrics, achieving an $R^{2}$ of 0.979 and an overall RMSE of 0.135 mm/day. Notably, the large model also accurately captures extreme precipitation events, with RMSE values of 0.365 mm/day at the 99.7th percentile and 0.525 mm/day at the 99.99th percentile. All RMSE values for precipitation are computed in log-transformed space using $\log(x+1)$, where $x$ denotes daily precipitation in millimeters.
Global Model Inference. Fig. 8 shows an animation of global daily precipitation for July 2020, comparing ORBIT-2 predictions (downscaling from 28 km to 7 km) against 7 km IMERG observations. The model was applied directly for global inference—without any fine-tuning or bias correction—to evaluate generalization. Despite challenges in downscaling ERA5 reanalysis data to satellite IMERG data due to their data source inconsistency, ORBIT-2 demonstrates strong generalization capability, achieving $R^{2}=0.90$, SSIM = 0.96, PSNR = 41.8, and RMSE = 0.34 mm/day (in $\log(x+1)$ space). These results underscore ORBIT-2’s robustness, effectively extending regional training to global inference.
VI Implication
@@
Crucially, ORBIT-2 achieves breakthrough scalability in ViT training, scaling up to 10 billion parameters model size across 32,768 GPUs, and scale up to 4.2 billion token sequence length, several magnitudes longer than the state-of-the-art long sequence implementation that scales to 188K tokens. It sustains up to 1.8 ExaFLOPs of performance with 92–98% strong scaling efficiency at 32,768 GPUs, setting a new benchmark for exascale AI workloads. ORBIT-2 serves as a blueprint for next-generation exascale foundation models, enabling transformative applications across domains including genomics, fluid dynamics, astrophysics, and Earth system modeling.
Impact on Earth System Science. For Earth system modeling, ORBIT-2 enables hyper-resolution, global-scale downscaling with state-of-the-art accuracy and efficiency. Evaluated against observational data, it achieves a coefficient of determination ($R^{2}$) of 0.999 for temperature and 0.979 for precipitation at 7 km resolution across the continental United States, closely matching ground-truth observation records. Critically, ORBIT-2 generalizes across both variables and geographic regions using a single unified foundation model—an essential capability for improving climate projection fidelity, detecting localized extremes, and guiding mitigation and adaptation strategies. This makes ORBIT-2 a valuable tool for researchers, policy-makers, and emergency planners worldwide.
@@ Acknowledgments
This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The U.S. Government retains a nonexclusive, worldwide license to publish or reproduce the published form of this manuscript, or to authorize others to do so, for U.S. Government purposes, as acknowledged by the publisher. This research was primary supported by the ORNL’s AI Initiative sponsored by the Director’s Research and Development Program at ORNL, additionally supported by DOE Early Career Project sponsored by the BER program. An award of computer time was provided by the INCITE program. This research also used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725.

### v2

@@ 7. Science Performance Results
U.S. Regional Fine-Tuning. Following pretraining on the datasets in Table 1, we fine-tune ORBIT-2 models with 9.5M and 126M parameters on two tasks for science demonstration: (1) US-specific downscaling of ERA5 and DAYMET data (1980 to 2022) from 28 km to 7 km, evaluated against 7 km DAYMET observations for both daily total precipitation and Tmin; and (2) global precipitation downscaling of ERA5 and IMERG (1998 to 2022) from 28 km to 7 km, evaluated against 7 km IMERG observations. These tasks evaluate model capability in downscaling at both regional and global scale.
@@
Table 6’s second row summarizes U.S. regional downscaling results for Tmin at the unseen testing years. Both 9.5M and 126M models accurately reconstruct high-resolution temperature at 7 km resolution and are capable of capturing extremes, with the larger 126M model consistently outperforming the smaller 9.5M model across all metrics with lower RMSE, and higher $R^{2}$, SSIM and PSNR. Notably, the 126M model achieves an $R^{2}$ of 0.999 and SSIM of 0.987, establishing a new benchmark for temperature downscaling at 7 km resolution.
Figure 8. Power spectra of downscaled minimum temperature using 9.5M and 126M parameter models, showing improved high-frequency fidelity with increased model capacity.
@@
Fig. 8 provides corresponding spectral analysis for Tmin downscaling results in Table 6’s row two, by comparing the spatial power spectra of the two model sizes. The 126M model accurately captures high-frequency content, closely matching the high frequency spectral characteristics of the DAYMET observation ground truth. In contrast, the 9.5M model deviates from the ground truth at high frequencies. This demonstrates the larger model’s ability to resolve fine-scale spatial variability, emphasizing the value of increased model capacity for high-fidelity climate downscaling.
@@
Table 6’s third row presents downscaling results for daily total precipitation—one of the most challenging variables due to its high spatial variability and localized extremes. ORBIT-2 demonstrates strong performance, closely matching observations when downscaling to 7 km resolution. The 126M model consistently outperforms the 9.5M model across all evaluation metrics, achieving an $R^{2}$ of 0.979 and an overall RMSE of 0.135 mm/day. Notably, the large model also accurately captures extreme precipitation events, with RMSE values of 0.365 mm/day at the 99.7th percentile and 0.525 mm/day at the 99.99th percentile. All RMSE values for precipitation are computed in log-transformed space using $\log(x+1)$, where $x$ denotes daily precipitation in millimeters.
Figure 9. Example region for global precipitation downscaling during fine-tuning. (a) ERA5 input at 28 km resolution; (b) ORBIT-2 downscaled output at 7 km. Click on the link here for an interactive visualization of global precipitation downscaling for the year 2020 Click Here.
Figure 10. (a) ERA5 skill scores for global precipitation at 28 km resolution, evaluated using averaged $R^{2}$ and SSIM, over 24 years (1998-2021). Analysis metrics include monsoon onset and withdrawal, seasonality, entropy, and Hovmöller diagnostics. (b) ORBIT-2 skill scores for global precipitation at 7 km resolution, showing improved fidelity across the same metrics.
Global Precipitation Fine-Tuning. ORBIT-2’s fine-tuning performance for global precipitation downscaling task is summarized in Table 6 row 4. Both model sizes (9.5M and 126M parameters) achieve similar accuracy, with $R^{2}=0.98$, SSIM = 0.92, PSNR = 35.0, and RMSE = 0.09 mm/day (in $\log(x+1)$ space). These results underscore ORBIT-2’s robustness across both regional and global downscaling tasks.
For visualization, Fig. 9 shows a one-day snapshot of an example region, comparing ERA5 precipitation at 28 km with ORBIT-2 downscaling at 7 km. A full-year animation is available online Click Here, illustrating the spatial resolution gains and improve climate pattern for year 2020.
To evaluate long-term fidelity and climate pattern, Fig. 10 reports performance of high-resolution downscaling in capturing key precipitation characteristics for 1998–2021 across 58 climatically homogeneous land and ocean regions defined by the Intergovernmental Panel on Climate Change (IPCC). For each region, we report a composite skill score, defined as the averaged $R^{2}$ correlation and SSIM (scaled between 0 to 100), with higher values indicating stronger agreement with the IMERG 7 km observational reference.
Performance is further analyzed in nine monsoon regions, where precipitation is governed by strong seasonality and complex land ocean atmosphere interactions, and significant hydrological and socio-economic impacts. Following (Ashfaq et al., 2021), we evaluate monsoon onset and withdrawal timing, precipitation seasonality and entropy, and the northward progression of rainfall (via Hovmöller diagnostics). Results show substantial improvements with ORBIT-2. As shown in Fig.10(a), ERA5 at 28 km resolution exhibits moderate skill scores across metrics and regions, whereas ORBIT-2 7 km downscaling (Fig.10(b)) consistently achieves much higher skill scores across all metrics and monsoon regions. These results highlight the effectiveness of ORBIT-2 in enhancing the spatiotemporal fidelity of precipitation, especially in regions governed by complex climatic processes.
8. Implication
@@
Crucially, ORBIT-2 achieves breakthrough scalability in ViT training, scaling up to 10 billion parameters model size across 65,536 GPUs, and scale up to 4.2 billion token sequence length, several magnitudes longer than the state-of-the-art long sequence implementation that scales to 188K tokens. It sustains up to 4.1 ExaFLOPs of performance with 74–98% strong scaling efficiency at 65,536 GPUs, setting a new benchmark for exascale AI workloads. ORBIT-2 serves as a blueprint for next-generation exascale foundation models, enabling transformative applications across domains including genomics, fluid dynamics, astrophysics, and Earth system modeling.
Impact on Climate Science. ORBIT-2 enables hyper-resolution, global-scale downscaling with state-of-the-art accuracy and efficiency. Against observations, it achieves an $R^{2}$ of 0.999 for temperature and 0.979/0.986 for precipitation at 7 km resolution over the continental United States and globally, respectively. Unlike traditional approaches, ORBIT-2 generalizes across variables and regions using a single foundation model, improving climate projection fidelity, resolving fine-scale processes, and detecting localized extremes. This unified capability makes ORBIT-2 a valuable tool for supporting climate adaptation and mitigation strategies. Besides the above capability, this paper also shows ORBIT-2 foundation AI model’s capability to perform global downscaling in 4 millisecond for 9.5M parameter model and 0.5 second for a 10B parameter model with only 8 GPUs through model inferencing, showing the AI model’s capability for near-real-time climate and weather prediction on edge devices. This is transformative compared to the conventional numerical simulation methods, which often takes hours and days to compute on a large supercomputer.
Limitations and Practical Considerations. Like all AI-based methods, ORBIT-2 may inherit biases from training data source and lacks strict enforcement of physical conservation laws, requiring caution in applications sensitive to energy or mass balance. Practical deployment will require careful consideration on application-based bias correction, physics-informed integration, and domain-specific fine-tuning to ensure reliability across diverse climates and regimes. Finally, while near-real-time inference on limited hardware is promising, robust evaluation in operational settings is needed before widespread deployment in climate services and decision-making.
@@ Acknowledgments
We thank Jonathan Coles for helping us get scalability numbers on the Alps cluster. This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The U.S. Government retains a nonexclusive, worldwide license to publish or reproduce the published form of this manuscript, or to authorize others to do so, for U.S. Government purposes, as acknowledged by the publisher. This research was primary supported by the ORNL’s AI Initiative sponsored by the Director’s Research and Development Program at ORNL, additionally supported by DOE Early Career Project sponsored by the BER program. It was also supported as part of the Energy Exascale Earth System Model (E3SM) project, funded by the U.S. Department of Energy Office of Science, Office of Biological and Environmental Research, and Earth Systems Model Development Program. An award of computer time was provided by the INCITE program through the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility.